OpenStudy (anonymous):

6^-1 minus 7^-1 (simplify)

jimthompson5910 (jim_thompson5910):

6^-1 minus 7^-1 1/6 - 1/7 7/42 - 6/42 1/42

OpenStudy (anonymous):

Come on that is easy

OpenStudy (anonymous):

maybe for you it is, I am frazzled and lost.

OpenStudy (anonymous):

Convert to fraction using negative exponent. Find lowest common multiple of the two and multiply numerator and denominator. Then subtract.

jimthompson5910 (jim_thompson5910):

6^-1 = 1/(6^1) = 1/6 So 6^-1 = 1/6 The same applies to 7^-1

jimthompson5910 (jim_thompson5910):

from there, I subtracted the fractions like normal

OpenStudy (anonymous):

Ok, thanks! I am overthinking it.